OpenStudy (anonymous):

Ignacio invested $2,748 in a savings account which earns 4.92% interest compounded quarterly. What will the account be worth in six years?

OpenStudy (anonymous):

\[P= P_0\left(1+ \frac r t\right)^{n t} \]

OpenStudy (anonymous):

\[ P=2748\left(1+ \frac {.0492}{4}\right)^{6 (4)} \]

OpenStudy (anonymous):

\[ P=2748\left(1+ \frac {.0492}{4}\right)^{6 (4)}=3685.01 \]

OpenStudy (anonymous):

A=P(1+I)N A=2748(1+4.92 DIVIDE BY 400)POWER OF 24 3685.01